OpenStudy (anonymous):

If x varies directly as y, and x = 48 when y = 16, find x when y = 5. How..? Thanks you!

OpenStudy (alexwee123):

use y = kx where k is a constant

OpenStudy (anonymous):

x = ky 48 =16k k =3 x =3y so when y =5 x = 3*5 x =??

OpenStudy (anonymous):

\[\large x = k \times y\] Find k by plugging x = 48 and y = 16 After finding k, put y = 5 and find x in the same equation..

OpenStudy (anonymous):